# Analyzing Variations in Dependency Distributions Due to Code Smell Interactions

### v3

@@ 1. Introduction
Modularity is a critical aspect of software maintainability. When excessive dependencies accumulate between modules, even small changes can trigger ripple effects across the codebase, making maintenance activities complex and costly (cerny2024maintainability, 1). Practitioners therefore stress reducing inter-module dependencies to maintain high modularity. However, our understanding of which characteristics of code drive dependencies remains limited. In this paper, we investigate a prominent suspect: code smells, examining whether the number and nature of dependencies change in the presence of multiple code smells.
Code smells are characteristics of source code that indicate potential design flaws (fowler1999refactoring, 2) and are believed to degrade maintainability (fowler1999refactoring, 2, 3, 4, 5, 6, 7). Interestingly, studies of isolated code smells often report non-significant impacts (sjoberg2012quantifying, 8, 9), while those on multiple code smells frequently find significant effects (singh2019reducing, 7, 10, 11). This pattern, along with other studies (pietrzak2006leveraging, 12, 13, 14, 15, 16), suggests that when multiple code smells are present, their interactions through static dependencies may play a more critical role in degrading maintainability, further complicating maintenance and refactoring. However, to the best of our knowledge, few studies empirically investigate the impact of such interactions on software quality (yamashita2013exploring, 15, 16), and none analyze their influence on dependency growth.
@@
  • We design and conduct a large-scale empirical analysis involving 116 Java systems and five types of code smells. All source code, data, and replication instructions are available in our replication package (blind_2025_15624918, 17).
  • We provide the first large-scale empirical evidence that code smell interactions are associated with an increasing number of dependencies, along with consistent change directions (increase or decrease) for certain dependency types. Our results offer actionable guidance for code smell prioritization, detection, and the development of more effective refactoring strategies, benefiting tool developers, practitioners, and researchers aiming to improve system maintainability by removing code smells and decoupling modules.
@@ 2.1. Code Smells
The term code smell was introduced by Kent Beck in the late 1990s (bansal2022categorical, 18) and gained popularity after Martin Fowler described 22 of them (fowler1999refactoring, 2). Code smells can be classified into method-level and class-level categories, as some affect the maintainability of a single method, while others influence the maintainability of an entire class (lanza2007object, 13).
In this study, we focus on the following five (lanza2007object, 13):
@@
We selected these smells because (i) they represent both class-level and method-level design flaws, (ii) they are among the most frequently considered and empirically investigated code smells in the software maintenance literature (kaur2020systematic, 19), (iii) they have often been empirically identified as among the most prevalent in codebases (reis2020code, 20, 21) and harmful to maintainability (li2007empirical, 22, 23, 24, 10), and (iv) Lanza and Marinescu (marinescu2005measurement, 25, 13) proposed a well-justified and explainable detection approach for them, called the metric-based approach, which uses a set of metrics with predefined thresholds to identify code smell symptoms. This approach has been adopted by many popular automatic detection tools, such as PMD [^1] https://pmd.github.io/.
@@ 2.2. Influence of Code Smells on Maintainability
Code smells are widely believed and have been empirically shown to compromise software maintainability (fowler1999refactoring, 2, 3, 26, 27, 28, 8, 4, 5, 29, 30, 31, 9, 6, 32, 33, 34, 35, 36, 37, 38, 39, 40, 10, 7, 41, 42, 43). A key observation is that code smells rarely exist in isolation; instead, multiple code smells often co-locate within the same class and collectively exacerbate maintainability issues (pietrzak2006leveraging, 12, 16, 44, 15, 45, 46, 47, 48, 11).
@@
Abbes et al. (abbes2011empirical, 44) investigated the effects of the Blob and Spaghetti code smells on code comprehensibility. They reported that classes exhibiting neither smell, or only one of the two, showed no measurable impact; however, when the two smells co-located in the same class, comprehension tasks required more time and effort and resulted in lower accuracy. Hamdi et al. (hamdi2021empirical, 47), focusing on Android applications, found that 51% of classes in 1,923 apps contained at least two smells, with 14 specific smell pairs frequently co-locating. Martins et al. (martins2021code, 46) and Santana et al. (santana2024exploratory, 48) both found that classes containing multiple types of code smells tend to exhibit higher complexity and lower cohesion than those with only one or none, with certain combinations being particularly detrimental, such as a God Class co-located with a Feature Envy. Furthermore, the developer survey by Martins et al. (martins2021code, 46) reported that code containing co-located smells was considerably harder to understand and modify. Santana et al. (santana2024unraveling, 11) additionally showed that classes with co-located code smells undergo significantly more frequent and extensive changes, indicating reduced code stability. These findings align with earlier studies (sjoberg2012quantifying, 8, 9), which reported that individual smells alone have limited effects on maintenance effort, implying that smell co-location is a more important factor in real maintenance tasks.
@@
Interestingly, it has been suggested that maintenance problems in large classes with co-located code smells are sometimes caused not by complexity itself, but by interactions between the smells through static dependencies (yamashita2012assessing, 49, 16). Building on this idea, Yamashita and Moonen (yamashita2013exploring, 15) conducted a study involving six professional developers who performed real change tasks in four Java systems. They recorded daily interviews and think-aloud sessions. The authors observed that code smells can interact with each other through static dependencies not only when co-located, but also across different classes. Such cases were referred to as coupling, where code smells in different classes are connected by static dependencies. Based on a qualitative analysis of developer feedback, they found that these interactions often hinder code comprehension and change propagation due to the ripple effects of static dependencies, thereby increasing maintenance effort and defect risk. To the best of our knowledge, this remains the only study that has examined the influence of such interactions on maintainability. Subsequent work has typically focused on identifying common patterns, such as which code smells tend to be collocated or coupled (yamashita2015inter, 16, 50, 51, 52, 45), but has rarely linked these findings to maintainability.
@@
In summary, previous research has extensively examined the influence of code smell co-location on maintainability, but has rarely considered the factor of interaction. The only study addressing this issue was qualitative and based on a small dataset (yamashita2013exploring, 15). Our work differs fundamentally by linking interaction to a maintainability metric, namely the number of static dependencies as an indicator of modularity. Using a significantly larger dataset, we apply statistical tests to enable quantitative assessment. Through a detailed analysis of each type of static dependency and its flow direction, we provide practical and actionable insights for improving modularity, guiding code smell refactoring, prioritization, and detection.
@@ 3.2. Dependency
where relation is the the abstract kind of dependency (e.g., call), and source artifact type and target artifact type are the leaf artifact types in the hierarchy, providing more detail and avoiding double counting (i.e., Class, Interface, Functional Method, Accessor, Constructor, and Field). Each dependency type is directional, flowing from the source artifact to the target artifact. We treat two dependency types as distinct whenever any of the three fields in the triple differ (for example, (call, FM, FM) $\neq$ (call, Constructor, FM)).
@@ 3.3. Interaction
Two artifacts, a1 and a2, interact if there exists a static dependency connecting them. The set of such dependencies constitutes the interaction set; if it is empty, no interaction occurs. Any interaction has the property of relative location, with a value of Same or Different, indicating whether the artifacts reside in the same or in different classes. This distinction is motivated by the observation that dependency distributions may vary with relative location (e.g., access data typically involves FM call an accessor when in different classes, but FM use a field when in the same class).
@@ 3.6. Data Analysis
The testing was conducted from two analytical perspectives: (i) Union: using the Union dataset to provide an overall overview and identify general trends; and (ii) Per system: testing each system’s data individually to determine whether code smell interactions vary across systems, i.e., whether the general trend holds within each system.
Hypotheses. For a code smell pair $(\mathrm{CS1},\mathrm{CS2})$, let $t_{1}$ and $t_{2}$ be interaction types, where $t_{1}=\mathrm{CS1\text{-}CS2}$ and $t_{2}\in$ {CS1-nonCS2,nonCS1-CS2}. Let $l\in\{\mathrm{Different},\mathrm{Same}\}$ denote the relative location, $d\in\{\mathrm{General\ analysis},\mathrm{Specific\ analysis}\}$ denote the dependent variable of interest, and $p\in\{\mathrm{Union},\mathrm{Per\ system}\}$ denote the analysis perspective. To answer RQ2, we test the following hypotheses:
@@
The Mann-Whitney U test (mann1947test, 53) was selected as the primary statistical test. This test was chosen because our data does not follow a normal distribution, and the two interaction types being compared are independent with unequal sample sizes (e.g., there are more interacting artifact pairs that are CS1-nonCS2 or nonCS1-CS2 than CS1-CS2, due to the larger number of non-code smells compared to code smells).
Since this test only determines whether a difference exists but does not measure its magnitude, we computed Cliff’s Delta ($\delta$), a non-parametric effect size measure that quantifies the strength of the observed difference. Cliff’s Delta can be positive or negative. A positive value indicates that code smell interactions are associated with an increased number of dependencies, while a negative value suggests a decrease. To interpret Cliff’s Delta, we adopted widely accepted thresholds (vargha2000critique, 54): negligible ($|\delta|<0.147$), small ($0.147\leq|\delta|<0.33$), medium ($0.33\leq|\delta|<0.474$), and large ($|\delta|\geq 0.474$). We refer to the sign of Cliff’s delta as the change direction, indicating whether it represents an increase or a decrease, and to the absolute value as the change magnitude, reflecting the strength of the difference.
Certain interaction types at specific relative locations for some code smell pairs are unlikely to occur due to the inherent nature of the code smells. For example, our dataset does not contain any class that is simultaneously a BC or GC and a DC, because BC and GC are complex classes, whereas DC are simple classes. Additionally, certain dependency types are impossible for specific flow directions within certain code smell pairs due to differences in granularity. For instance, a (use, FM, Field) dependency cannot occur from GC to BM, as BM is a method-level smell and therefore cannot serve as a field type target artifact. Furthermore, some projects may not contain certain code smells, or may contain them without any interactions. In such cases, statistical tests cannot be performed. However, the conclusions of this study remain valid, as we focus only on cases where interactions occur.
@@ 3.8. Implementation
CodeQL [^3] https://codeql.github.com/ is a static analysis tool originally developed to identify vulnerabilities in codebases. Beyond security analysis, CodeQL allows users to detect custom code patterns using a declarative query language (youn2023declarative, 55). In this study, it was employed to detect code smells, analyze dependencies, and examine interactions. We employed the popular and widely used metric-based approach proposed by Lanza and Marinescu (lanza2007object, 13) to detect code smells, as described in Section 2.1. Our detection queries were implemented using the metrics and thresholds they provided (see Section 5.5 for why we built our own smell detector). When a threshold was specified as a range, we selected the stricter bound to minimize false positives. That is, for metrics required to be smaller than the threshold, we selected the lower bound, and for metrics required to be larger than the threshold, we selected the upper bound. For each project, we excluded third-party libraries from analysis, as developers typically do not have control over their source code. Our ultimate goal is to ensure that our findings are actionable for developers. Additionally, we excluded test and example code and focused exclusively on production code. See our replication package for details on code smell and metric definitions, as well as the chosen thresholds and detection rules (blind_2025_15624918, 17).
@@
Evaluation. To validate detection correctness, we computed the per-smell sample size using the standard confidence interval for a proportion (95% confidence, $\pm 10\%$ margin) and then performed stratified proportional sampling by project: for each smell, we allocated $n_{i}\propto N_{i}$ (where $N_{i}$ is the number of detections in project $i$) and drew simple random samples within each project (cochran1977sampling, 56). This yields an unbiased lower-variance estimate of the overall false positive rate while ensuring representation is aligned with project contribution (singh1996stratified, 57). The first author then performed a manual evaluation of the sampled smell instances, first checking if the metric measurement was correct, and then checking if it was strictly aligned with the definition of that smell provided by Lanza and Marinescu (lanza2007object, 13). Our tool also outputs information related to each metric (e.g., for metrics used to detect FE, it reports which foreign fields are accessed and from which classes.), which allows us to quickly verify the measurement by checking if that information is correct. At the end, we manually checked 94, 94, 93, 86, and 95 samples for FE, BM, GC, BC, and DC, respectively. All sampled detections and their metrics were judged correct (100%), with no false positives observed for any smell, indicating that the detected smell instances are consistent with the definitions we used.

### v4

# Analyzing Dependency Distribution Changes Arising from Code Smell Interactions

@@ Abstract.
code smells, static analysis, static dependencies, modularity, code smell interactions, code smell co-occurrence, code smell agglomeration, software maintenance
journalyear: 2026copyright: ccconference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793302.3793365isbn: 979-8-4007-2474-9/2026/04ccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software post-development issuesccs: Software and its engineering Abstraction, modeling and modularityccs: Software and its engineering Software evolution
@@ 1. Introduction
Modularity is a critical aspect of software maintainability. When excessive dependencies accumulate between modules, even small changes can trigger ripple effects across the codebase, making maintenance activities complex and costly (Cerny et al., 2024). Practitioners therefore stress reducing inter-module dependencies to maintain high modularity. However, our understanding of which characteristics of code drive dependencies remains limited. In this paper, we investigate a prominent suspect: code smells, examining whether the number and nature of dependencies change in the presence of multiple code smells.
Code smells are characteristics of source code that indicate potential design flaws (Fowler, 1999) and are believed to degrade maintainability (Fowler, 1999; Munro, 2005; Marinescu, 2012; Yamashita and Counsell, 2013; Soh et al., 2016; Singh et al., 2019). Interestingly, studies of isolated code smells often report non-significant impacts (Sjøberg et al., 2013; Yamashita, 2014), while those on multiple code smells frequently find significant effects (Singh et al., 2019; Palomba et al., 2018; Santana et al., 2024b). This pattern, along with other studies (Pietrzak and Walter, 2006; Lanza and Marinescu, 2007; Moha et al., 2009; Yamashita and Moonen, 2013; Yamashita et al., 2015), suggests that when multiple code smells are present, their interactions through static dependencies may play a more critical role in degrading maintainability, further complicating maintenance and refactoring. However, to the best of our knowledge, few studies empirically investigate the impact of such interactions on software quality (Yamashita and Moonen, 2013; Yamashita et al., 2015), and none analyze their influence on dependency growth.
@@
  • We designed and conducted a large-scale empirical analysis involving 116 Java systems and five types of code smells. All source code, data (e.g., code smell detection output, metric measurements for code smell detection, dependency detection output, and statistical analysis results), and replication instructions are available in our replication package (Zhang et al., 2025).
  • We provide the first large-scale empirical evidence that code smell interactions are associated with an increasing number of dependencies, along with consistent change directions (increase or decrease) for certain dependency types. Our results offer actionable guidance for code smell prioritization and detection, as well as for developing more effective refactoring strategies, benefiting tool developers and practitioners aiming to improve software maintainability by removing code smells and decoupling modules. Researchers can also build on our study design and findings on changes in dependency distribution to further investigate code smell interactions.
@@ 2.1. Code Smells
The term code smell was introduced by Kent Beck in the late 1990s (Bansal et al., 2022) and gained popularity after Martin Fowler described 22 of them (Fowler, 1999). Code smells can be classified into method-level and class-level categories, as some affect the maintainability of a single method, while others influence the maintainability of an entire class (Lanza and Marinescu, 2007).
In this study, we focus on the following five (Lanza and Marinescu, 2007):
@@
We selected these smells because (i) they represent both class-level and method-level design flaws, (ii) they are among the most frequently considered and empirically investigated code smells in the software maintenance literature (Kaur, 2020), (iii) they have often been empirically identified as among the most prevalent in codebases (Reis et al., 2020; Zakeri-Nasrabadi et al., 2023) and harmful to maintainability (Li and Shatnawi, 2007; Cairo et al., 2018; Tufano et al., 2017; Palomba et al., 2018), and (iv) Lanza and Marinescu (Marinescu, 2005; Lanza and Marinescu, 2007) proposed a well-justified and explainable detection approach for them, called the metric-based approach, which uses a set of metrics with predefined thresholds to identify code smell symptoms. This approach has been adopted by many popular automatic detection tools, such as PMD [^1] https://pmd.github.io/.
@@ 2.2. Influence of Code Smells on Maintainability
Code smells are widely believed and have been empirically shown to compromise software maintainability (Fowler, 1999; Munro, 2005; Lozano and Wermelinger, 2008; Olbrich et al., 2010; D’Ambros et al., 2010; Sjøberg et al., 2013; Marinescu, 2012; Yamashita and Counsell, 2013; Sabané et al., 2013; Chatterji et al., 2013; Bán and Ferenc, 2014; Yamashita, 2014; Soh et al., 2016; Rahman and Roy, 2017; Mondal et al., 2017; Zhang et al., 2017; Islam et al., 2017; Bán, 2017; Jaafar et al., 2017; Elish, 2017; Mondal et al., 2018; Barbour et al., 2018; Palomba et al., 2018; Singh et al., 2019; Aversano et al., 2020; Imran and Kosar, 2022; Masudur Rahman et al., 2025). A key observation is that code smells rarely exist in isolation; instead, multiple code smells often co-locate within the same class and collectively exacerbate maintainability issues (Pietrzak and Walter, 2006; Yamashita et al., 2015; Abbes et al., 2011; Yamashita and Moonen, 2013; Martins et al., 2020, 2021; Hamdi et al., 2021; Santana et al., 2024a, b).
@@
Abbes et al. (Abbes et al., 2011) investigated the effects of the Blob and Spaghetti code smells on code comprehensibility. They reported that classes exhibiting neither smell, or only one of the two, showed no measurable impact; however, when the two smells co-located in the same class, comprehension tasks required more time and effort and resulted in lower accuracy. Hamdi et al. (Hamdi et al., 2021), focusing on Android applications, found that 51% of classes in 1,923 apps contained at least two smells, with 14 specific smell pairs frequently co-locating. Martins et al. (Martins et al., 2021) and Santana et al. (Santana et al., 2024a) both found that classes containing multiple types of code smells tend to exhibit higher complexity and lower cohesion than those with only one or none, with certain combinations being particularly detrimental, such as a God Class co-located with a Feature Envy. Furthermore, the developer survey by Martins et al. (Martins et al., 2021) reported that code containing co-located smells was considerably harder to understand and modify. Santana et al. (Santana et al., 2024b) additionally showed that classes with co-located code smells undergo significantly more frequent and extensive changes, indicating reduced code stability. These findings align with earlier studies (Sjøberg et al., 2013; Yamashita, 2014), which reported that individual smells alone have limited effects on maintenance effort, implying that smell co-location is a more important factor in real maintenance tasks.
@@
Interestingly, it has been suggested that maintenance problems in large classes with co-located code smells are sometimes caused not by complexity itself, but by interactions between the smells through static dependencies (Yamashita, 2012; Yamashita et al., 2015). Building on this idea, Yamashita and Moonen (Yamashita and Moonen, 2013) conducted a study involving six professional developers who performed real change tasks in four Java systems. They recorded daily interviews and think-aloud sessions. The authors observed that code smells can interact with each other through static dependencies not only when co-located, but also across different classes. Such cases were referred to as coupling, where code smells in different classes are connected by static dependencies. Based on a qualitative analysis of developer feedback, they found that these interactions often hinder code comprehension and change propagation due to the ripple effects of static dependencies, thereby increasing maintenance effort and defect risk. To the best of our knowledge, this remains the only study that has examined the influence of such interactions on maintainability. Subsequent work has typically focused on identifying common patterns, such as which code smells tend to be collocated or coupled (Yamashita et al., 2015; Fontana et al., 2015; Walter et al., 2018; Sobrinho and Maia, 2021; Martins et al., 2020), but has rarely linked these findings to maintainability.
@@
In summary, previous research has extensively examined the influence of code smell co-location on maintainability, but has rarely considered the factor of interaction. The only study addressing this issue was qualitative and based on a small dataset (Yamashita and Moonen, 2013). Our work differs fundamentally by linking interaction to a maintainability metric, namely the number of static dependencies as an indicator of modularity. Using a significantly larger dataset, we apply statistical tests to enable quantitative assessment. Through a detailed analysis of each type of static dependency and its flow direction, we provide practical and actionable insights for improving modularity, guiding code smell refactoring, prioritization, and detection.
@@ 3.2. Dependency
where relation is the abstract kind of dependency (e.g., call), and source artifact type and target artifact type are the leaf artifact types in the hierarchy, providing more detail and avoiding double-counting (i.e., Class, Interface, Functional Method, Accessor, Constructor, and Field). Each dependency type is directional, flowing from the source artifact to the target artifact. We treat two dependency types as distinct whenever any of the three fields in the triple differ (for example, (call, FM, FM) $\neq$ (call, Constructor, FM)).
@@ 3.3. Interaction
Interaction Example
Two artifacts, a1 and a2, interact if there exists a static dependency connecting them. The set of such dependencies constitutes the interaction set; if it is empty, no interaction occurs. Any interaction has the property of relative location, with a value of Same or Different, indicating whether the artifacts reside in the same or in different classes. This distinction is motivated by the observation that dependency distributions may vary with relative location (e.g., when accessing data in different classes, FM typically call accessors, but when in the same class, FM typically use fields).
@@ 3.6. Data Analysis
The testing was conducted from two analytical perspectives: (i) Union: using the Union dataset to provide an overall overview and identify general trends; and (ii) Per-system: testing each system’s data individually to determine whether code smell interactions vary across systems, i.e., whether the general trend holds within each system.
Hypotheses. For a code smell pair $(\mathrm{CS1},\mathrm{CS2})$, let $t_{1}$ and $t_{2}$ be interaction types, where $t_{1}=\mathrm{CS1\text{-}CS2}$ and $t_{2}\in$ {CS1-nonCS2,nonCS1-CS2}. Let $l\in\{\mathrm{Different},\mathrm{Same}\}$ denote the relative location, $d\in\{\mathrm{General\ analysis},\mathrm{Specific\ analysis}\}$ denote the dependent variable of interest, and $p\in\{\mathrm{Union},\text{Per-system}\}$ denote the analysis perspective. To answer RQ2, we test the following hypotheses:
@@
The Mann-Whitney U test (Mann and Whitney, 1947) was selected as the primary statistical test. This test was chosen because our data does not follow a normal distribution, and the two interaction types being compared are independent with unequal sample sizes (e.g., there are more interacting artifact pairs that are CS1-nonCS2 or nonCS1-CS2 than CS1-CS2, due to the larger number of non-code smells compared to code smells).
Since this test only determines whether a difference exists but does not measure its magnitude, we computed Cliff’s Delta ($\delta$), a non-parametric effect size measure that quantifies the strength of the observed difference. Cliff’s Delta can be positive or negative. A positive value indicates that code smell interactions are associated with an increased number of dependencies, while a negative value suggests a decrease. To interpret Cliff’s Delta, we adopted widely accepted thresholds (Vargha and Delaney, 2000): negligible ($|\delta|<0.147$), small ($0.147\leq|\delta|<0.33$), medium ($0.33\leq|\delta|<0.474$), and large ($|\delta|\geq 0.474$). We refer to the sign of Cliff’s Delta as the change direction, indicating whether it represents an increase or a decrease, and to the absolute value as the change magnitude, reflecting the strength of the difference.
Certain interaction types at specific relative locations for some code smell pairs are unlikely to occur due to the inherent nature of the code smells. For example, our dataset does not contain any class that is simultaneously a BC or GC and a DC, because BC and GC are complex classes, whereas DC are simple classes. Additionally, certain dependency types are impossible for specific flow directions within certain code smell pairs due to differences in granularity. For instance, a (use, FM, Field) dependency cannot occur from GC to BM, as BM is a method-level smell and therefore cannot serve as a field type target artifact. Furthermore, some projects may not contain certain code smells or may contain them without any interactions. In such cases, statistical tests cannot be performed. However, the conclusions of this study remain valid, as we focus only on cases where interactions occur.
Multiple Testing Correction. To reduce the risk of false discoveries, we applied the Benjamini-Hochberg procedure in our analyses. For the Union dataset, we adjusted across 10 (pairs) $\times$ 4 (contrasts) =40=40 p-values for the general analysis, and 10 $\times$ 4 $\times$ 2 (flow directions) $\times$ 31 (dependency types; considering all possible triples in Table 1) =2,480=2{,}480 p-values for the specific analysis. For the per-system dataset, we adjusted across $40\times 116$ (number of systems)=4,640=4{,}640 p-values for the general analysis, and $2{,}480\times 116=287{,}680$ p-values for the specific analysis. Note that these numbers represent upper bounds; the actual number of p-values is lower because some cases are unlikely to occur, as explained in the previous paragraph.
@@ 3.8. Implementation
CodeQL [^3] https://codeql.github.com/ is a static analysis tool originally developed to identify vulnerabilities in codebases. Beyond security analysis, CodeQL allows users to detect custom code patterns using a declarative query language (Youn et al., 2023). In this study, it was employed to detect code smells, analyze dependencies, and examine interactions. We employed the popular and widely used metric-based approach proposed by Lanza and Marinescu (Lanza and Marinescu, 2007) to detect code smells, as described in Section 2.1. Our detection queries were implemented using the metrics and thresholds they provided (see Section 5.5 for why we built our own smell detector). When a threshold was specified as a range, we selected the stricter bound to minimize the false positive rate. That is, for metrics required to be smaller than the threshold, we selected the lower bound, and for metrics required to be larger than the threshold, we selected the upper bound. For each project, we excluded third-party libraries from analysis, as developers typically do not have control over their source code. Our ultimate goal is to ensure that our findings are actionable for developers. Additionally, we excluded test and example code and focused exclusively on production code. See our replication package for details on code smell and metric definitions, as well as the chosen thresholds and detection rules (Zhang et al., 2025).
@@
Evaluation. To validate detection correctness, we computed the per-smell sample size using the standard confidence interval for a proportion (95% confidence, $\pm 10\%$ margin) and then performed stratified proportional sampling by project: for each smell, we allocated $n_{i}\propto N_{i}$ (where $N_{i}$ is the number of detections in project $i$) and drew simple random samples within each project (Cochran, 1977). This yields an unbiased, lower-variance estimate of the overall false positive rate while ensuring that representation is aligned with each project’s contribution (Singh and Mangat, 1996). The first author then performed a manual evaluation of the sampled smell instances, first checking whether the metric measurements were correct and then checking whether each detected instance was strictly aligned with the definition of that smell provided by Lanza and Marinescu (Lanza and Marinescu, 2007). Our tool also outputs information related to each metric (e.g., for metrics used to detect FE, it reports which foreign fields are accessed and from which classes), which allows us to quickly verify the measurements by checking whether this information is correct. In total, we manually checked 94, 94, 93, 86, and 95 samples for FE, BM, GC, BC, and DC, respectively. All sampled detected smell instances and their metric measurements were judged correct (100%), with no false positives observed in the sample, indicating that these instances are consistent with the definitions we used. We also applied the same procedure to non-code smells by manually checking 97, 97, 96, 96, and 96 samples for nonFE, nonBM, nonGC, nonBC, and nonDC, respectively. All metric measurements were judged correct, and the false negative rates were low: 2.1%, 6.2%, 9.3%, 8.3%, and 11.5%, respectively.
Replication. All analyses were executed on a desktop machine (Windows 10, AMD Ryzen 9 5950X, 32 GB RAM). To replicate the results, run the provided CodeQL queries on the pre-built CodeQL databases and execute the accompanying Python scripts. A Java runtime is required only to build CodeQL databases; because we provide pre-built databases, Java is not required to replicate our results. On our machine, the full pipeline completed in approximately 5 hours. Runtime may vary with CPU performance (e.g., core count and clock speed) and storage performance. Detailed replication instructions are included in the replication package (Zhang et al., 2025).